# Widespread disruption of resonant chains during protoplanetary disk dispersal

### v1

@@ Abstract
We present simulations of three planet systems with masses based on the observed triple planet systems observed by the Kepler satellite within the context of this model. The final distribution of nearest neighbour period ratios provides an excellent fit to the observations, provided that the initial resonant configurations are compact. The occurrence rate of planets as a function of orbital period also provides a good match to the observations, for final orbital periods $P<20$ days.
@@ 1 Introduction
For models in which the planets migrate within the protoplanetary disk, a critical element is the process that determines the halting of migration. In particular, the structure of the inner edge of the disk is expected to be regulated by the interaction with the stellar magnetic field (Romanova & Lovelace, 2006; Romanova et al., 2019) and the change in the density structure reduces or reverses the torque driving the planets inwards (Masset et al., 2006; Tsang, 2011; Miranda & Lai, 2018). For the purposes of modelling migration, this is often modelled as a simple truncation, resulting from the balance of the ram pressure of the disk flow with the magnetic pressure of the magnetic field (Terquem & Papaloizou, 2007; Chang et al., 2010; Izidoro et al., 2017; Brasser et al., 2018). However, Yu et al. (2023) have shown that a more accurate model – accounting for the diffusion of the magnetic field into the disk – results in an inner transition region, where the torque on the planet is not zero but changes sign, driving outwards migration. Thus, migrating planets reverse the course of their migration as they cross this transition and therefore pile up not at the edge of the magnetosperic cavity, but at a more distant location. As the accretion rate drops, the torque reversal location moves outwards, causing the planet to move outwards again, coming to rest at the point where the disk density has dropped to the point that the migration ‘freezes out’.
In Yu et al. (2023) we demonstrated how this model determines the final orbital periods in the case of a single planet. In this paper, we wish to investigate the consequences of this model for multiple planet systems. An important feature of the multiple planet case is that it can result in planetary pairs breaking out of resonance. Liu et al. (2017) have shown that, as the accretion rate drops, the magnetospheric cavity expands and can drive outward migration of a planet trapped at the inner edge. They use a one-sided torque operating at the disk edge, based on the results of the studies above. The fact that outward torques operate only on one planet limits the degree to which this effect operates to disrupt a resonance chain, and the resulting simulations still leave too many systems in resonant configurations (Liu & Ormel, 2017). As we will show, the presence of a non-zero torque interior to the location of the torque reversal results in multiple planets migrating outwards and provides a more extensive dynamical perturbation to the system.
@@ 2 Dynamical Model
We wish to model the resonant dynamics of two and three planet systems, under the influence of the most important resonant interactions. We choose to focus on the first order resonances $(q+1)/q$ where $q=1$ and $q=2$. We will also include the effects of torques from the protoplanetary disk on both semi-major axis evolution and planetary eccentricity damping. This is done by including, in the equations, a linear decay term characterised by timescale, as has been done by several prior authors – (Terquem & Papaloizou, 2019; Choksi & Chiang, 2020). One difference in our case is that our disk model is based on the model of Yu et al. (2023), which features a reversal of the torque direction in the inner disk, due to the penetration of the stellar magnetic field into the disk. As a result, the sign of the torque can reverse, so that we must allow for the timescale for semi-major axis evolution ($\tau_{a}$) to become negative. The timescale for eccentricity damping is always positive. The equations featuring three planets and two resonances are sufficiently lengthy that we defer their explicit expression to appendix A.
@@ 2.1 Torque Model
The dynamical evolution of the system is driven by the torques exerted by the protoplanetary disks. We determine a functional form for the torque based on a fit to the numerical disk simulations discussed in Yu et al. (2023). For the inward migration, the characteristic timescale for semi-major axis evolution is
$\tau_{a}$ $=$ $0.36{\rm Myr}\left(\frac{a}{0.1AU}\right)^{1.38}\left(\frac{M_{p}}{10M_{\oplus}}\right)^{-1}$ (1)
@@
The time dependance of $\tau_{a}$ is driven by the evolution of the accretion rate, which is
@@
This is chosen to represent a disk that decays exponentially as it accretes onto the star but then cuts off more rapidly once photoevaporation losses detach the inner disk from mass resupply by the outer disk (Alexander et al., 2014).
Once a planet has crossed interior to the torque reversal line $a_{R}$, the migration is now directed outwards. We describe this using a negative timescale. The functional form is again determined by fitting to the simulations of Yu et al. (2023), and is expressed as a function of position relative to $a_{R}$, $z=\ln a_{1}-a_{R}$. This amounts to
$\tau_{a}$ $=$ $6.66\times 10^{-18}{\rm Myr}\left(\frac{0.1AU}{a}\right)^{1/2}\left(\frac{10M_{\oplus}}{M_{p}}\right)$
@@
in the case of our default model. Indeed, if the migration time inward or outward is shorter than the timescale on which $a_{R}$ is evolving, then the planet will track the torque reversal location, making the relevant characteristic timescale $\tau_{R}=dt/d\ln a_{R}$.
@@
For example, if the outer planet of the pair reaches the torque reversal location, then the pair of planets will both be migrating outwards. The question of whether they converge or diverge will rest on whether the outward migration of the inner planet is faster or slower than the rate at which $a_{R}$ recedes. Two examples of this case are shown in Figure 1. The solid curve shows the characteristic timescale associated with the outward movement of the torque reversal location as $\dot{M}$ decreases. We also show two cases of $\tau_{a}$ interior to this. The dotted line shows the rate of outward migration for a $6M_{\oplus}$ planet that is the interior member of a pair locked in the 2:1 mean motion resonance, in which the outer member is located at $a_{R}$. We see that this increases with time, while the solid curve decreases with time. Thus, initially, the inner planet is moving outwards faster, and remains locked in resonance because the pair is converging. However, when the two curves cross, the inner planet can no longer keep up and the pair starts to diverge.
@@
Figure 1: The solid curve shows the characteristic timescale for the evolution of $a_{R}$ as a function of time. We see that this is initially much longer than the lifetime of the disk, but shortens as time increases, i.e. the torque reversal moves out faster as $\dot{M}$ drops. The dotted curve represents $\tau_{a}$ at a location $0.63a_{R}$, assuming a mass $6M_{\oplus}$. This represents the rate of outward migration for a planet of that mass, trapped in a 2:1 resonance with a planet located at the torque reversal location. The dashed line represents the same but for a $3M_{\oplus}$ planet located at $0.763a_{R}$, i.e. in a 3:2 resonance. The long dashed line indicates the timescale for inward migration, evaluated at $a_{R}$. Both timescales are shorter than $\tau_{R}$ until about 1.2 Myr. After this, the planets find it harder to track $a_{R}$ and will freeze out and decouple from the disk.
@@
The dashed line shows the case of a $3M_{\oplus}$ planet locked in a 3:2 resonance with an outer member located at $a_{R}$. This shows the same characteristic behaviour, although the divergence starts earlier.
@@ 3 Converging versus Diverging
where we have chosen the q=1 resonance for this example and the various quantities are defined in appendix A. These conditions also imply specific values for the equilibrium eccentricities. The full expressions are given in Terquem & Papaloizou (2019), but the most important element is that
@@
Thus, when the innermost planet first crosses the torque reversal location, the value of $t_{a,1}$ will change sign. This will cause the pair to converge even more strongly, so that there will still be an equilibrium, although the values of the equilibrium eccentricities will change. However, once the second planet crosses the torque reversal location, then both planets are moving outwards and the question of whether they are converging or diverging is once again a matter of their relative motion. The equilibrium will fail if the quantity in equation (6) becomes negative – thereby breaking the resonant lock. The system evolution retains the same character if one of the planets is tracking the location of the torque reversal – the relative $t_{a}$ is then simply dictated by the evolution of the torque reversal.
@@ 4 Two Planet Systems
Figure 2 illustrates the basic evolution for a two planet system, containing two $3M_{\oplus}$ planets. The black curve in the upper left plot illustrates the evolution of the semi-major axis of the inner planet, and the red curve the evolution of the outer planet. The dotted line indicates the evolution of the location of the torque reversal location $a_{R}$. The upper right panel shows the period ratio of the two planets as a function of time. In the lower left, we show the evolution of the planetary eccentricities, and, in the lower right, the panel shows how the resonant angle for the inner and outer 3:2 resonances change with the planetary and disk evolution.
@@
Figure 2: The upper left panel shows the evolution of semi-major axis for a pair of $3M_{\oplus}$ planets migrating due to the gaseous torque model described in the text. The dotted curve shows how the location of the torque reversal evolves with time. Important transitions in the dynamical state of the system are labelled as A–E, and are show as vertical dashed lines in the other three panels. In the upper right, we see the evolution of the period ratio for the pair. In the lower left we show the evolution of the eccentricity for each planet (red curve indicating the outer planet and black curve the inner planet). In the bottom right, we see the evolution of the resonant angles $3\lambda_{2}-2\lambda_{1}-\bar{\omega}_{1}$ (black) and $3\lambda_{2}-2\lambda_{1}-\bar{\omega}_{2}$ (red).
@@
The first incident of note – denoted point A – is when the inner planet reaches the torque reversal location $a_{R}$. At this point the inward migration of the inner planet stops and the planet semi-major axis starts to track the evolution of $a_{R}$, which is slowly moving outwards. Up until this point, the period ratio is increasing (the planets are diverging) because the migration time decreases as planets get closer to the star. However, with the halting of the inner planet migration at A, the planets begin to converge because the outer planet, shown as the red curve, continues to migrate inwards, and the two planets enter the 3:2 resonance at B. The net relative motion is convergent and so the pair become locked in a resonant configuration. They also acquire finite equilibrium eccentricities, as expected from the discussion in § 3.
@@
The torques on the outer planet now drive the pair inwards, and the inner planet is pushed interior to the torque reversal. Eventually the outer planet reaches the torque reversal (at time denoted as C). The planets remain in equilibrium (because the torques on the inner planet are pushing it outwards, so the pair is still converging) but the equiliibrium eccentricities change because the sign of the torques change. After this, the pair starts to migrate outwards but remains in resonance as long as the rate of outward migration of the inner planet is faster than the evolution of the torque reversal. This continues for a long time, but eventually, as the torques weaken and the outward motion of $a_{R}$ accelerates, the pair starts to diverge, at epoch D. The libration in the 3:2 resonance begins to increase, as does the period ratio, and the eccentricities start to damp away as their resonant excitation weakens. Eventually the migration freezes out completely (at E), leaving the system still in resonance with a finite libration amplitude and an offset relative to strict commensurability.
@@
This is the fundamental result that emerges from these calculations – in the disk models of Yu et al. (2023), planets are not trapped at the true inner edge of the disk, but rather at the location of the torque reversal. As the disk evolves and the surface density drops, this location moves outwards, resulting in an outward migration of the planets at, and interior to, the torque reversal. In many cases, this leads to divergent migration, thereby breaking the resonant lock and causing divergence of the period ratios away from the initial commensurabilities, to values more consistent with the observations.
@@
The planetary masses also play a role in determining the outcome of the evolution, because more massive planets remain coupled to the gas disk for longer, and lower mass planets freeze out earlier. To illustrate this, Figure 3 shows the case of a two planet system wherein the outer planet ($6M_{\oplus}$) is more massive than the inner one ($3M_{\oplus}$). The initial evolution is qualitatively similar, except that the stronger torques on the outer planet means that the inner planet barely halts at the torque reversal before being captured in resonance and pushed further inwards. The eventual divergence from the 3:2 resonance (at D) is also qualitatively similar. However, the outer planet remains coupled to the gas disk for longer by virtue of its larger mass and so migrates far enough that the system passes through the 2:1 resonance (shown at D). This resonant encounter is now divergent, so there is no capture (the system leaves the 2:1 resonance at E) but there is a transient excitation of the eccentricities that occurs late enough to leave small non-zero eccentriities ($\sim 0.01$) after the system has frozen out (at F).
@@
Figure 3: The upper left panel shows the evolution of semi-major axis for a pair of migrating planets where the outer planet is $6M_{\oplus}$ and the inner planet is $3M_{\oplus}$. The dotted curve shows how $a_{R}$ evolves with time. Important transitions in the dynamical state of the system are labelled as A–F, and are shown as vertical dashed lines in the other three panels. In the upper right, we see the evolution of the period ratio for the pair. In the lower left we show the evolution of the eccentricity for each planet (red curve indicating the outer planet and black curve the inner planet). In the bottom right, we have two panels to track the evolution of the resonant angles for the two important first order resonances that are encountered. The one designated as $\phi_{3:2}$ shows the angles $3\lambda_{2}-2\lambda_{1}-2\bar{\omega}_{1}$ (black) and $3\lambda_{2}-2\lambda_{1}-2\bar{\omega}_{2}$ (red). The one designated as $\phi_{2:1}$ shows the evolution of the two resonant angles $2\lambda_{2}-\lambda_{1}-\bar{\omega}_{1}$ (black) and $2\lambda_{2}-\lambda_{1}-\bar{\omega}_{2}$ (red).
@@
The overall amount of migration is driven largely by the planetary masses, as they determine how long the planet remains coupled to the disk and how rapidly the planets migrate with respect to one another. Figure 4 shows the effect of both planetary mass ratio and total mass in the overall evolution of a range of planet pairs. We see that low mass ratios ($M_{2}/M_{1}<0.6$) leave pairs trapped in resonance, whereas larger mass ratios lead to divergence and a wide range of final period ratios. This is a consequence of the fact that the more massive planets remain coupled to the gas for longer. Thus, if the inner planet is the more massive one, the pair will remain in the convergent state for the entirety of the disk evolution. However, if the outer planet is the more massive one, then the inner planet will freeze out first and fall behind the outer planet’s migration. This leads to a diverging configuration and a final period ratio that increases as $M_{2}/M_{1}$ does. The degree of divergence also depends on the total planet mass, with more massive planetary systems leading to large final mass ratios (because they remain coupled to the disk for longer).
@@
Figure 4: The lower panel shows the final period ratio of a planetary pair as a function of the mass ratio $M_{2}/M_{1}$. The results are color coded according to the final value, with black signifying $P_{2}/P_{1}<1.55$, red (1.55–1.95), magenta (1.95–2.05), green (2.05–3) and cyan for $P_{2}/P_{1}>3$. In the upper panel, these colors indicate the final states of systems as a function of mass ratio and total system mass $M_{1}+M_{2}$. The dotted lines in the lower panel indicate the first order commensurabilities (1+q)/q for q=2 and q=1.

### v2

@@ Abstract
We present simulations of three planet systems, with masses based on the observed triple planet systems observed by the Kepler satellite, within the context of this model. The final distribution of nearest neighbour period ratios provides an excellent fit to the observations, provided that the initial configurations contain a significant fraction of pairs whose period ratios are less than 2:1. The occurrence rate of planets as a function of orbital period also provides a good match to the observations, for final orbital periods $P<20$ days.
@@ 1 Introduction
For models in which the planets migrate within the protoplanetary disk, a critical element is the process that determines the halting of migration. In particular, the structure of the inner edge of the disk is expected to be regulated by the interaction with the stellar magnetic field (Romanova & Lovelace, 2006; Romanova et al., 2019) and the change in the density structure reduces or reverses the torque driving the planets inwards (Masset et al., 2006; Tsang, 2011; Miranda & Lai, 2018). For the purposes of modelling migration, this is often presented as a simple truncation in surface density, resulting from the balance of the ram pressure of the disk flow with the magnetic pressure of the magnetic field (Terquem & Papaloizou, 2007; Chang et al., 2010; Izidoro et al., 2017; Brasser et al., 2018). However, Yu et al. (2023) have shown that a more accurate model, that accounts for the diffusion of the magnetic field into the disk, will result in a surface density that increases outwards from the inner edge before reaching a maximum and reverting to a more traditional power-law decay to larger radii. The change in the surface density and temperature profiles reverses the torque direction interior to the surface density maximum. Thus, migrating planets reverse the course of their migration as they cross this transition and pile up, not at the edge of the magnetosperic cavity, but at a more distant location – the surface density maximum. As the accretion rate drops, the torque reversal location moves outwards, causing the planet to move outwards again, coming to rest at the point where the disk density has dropped to the point that the migration ‘freezes out’.
In Yu et al. (2023) we demonstrated how this model determines the final orbital periods in the case of a single planet. In this paper, we wish to investigate the consequences of this model for multiple planet systems. An important feature of the multiple planet case is that it can result in planetary pairs breaking out of resonance. Liu et al. (2017) have shown that, as the accretion rate drops, the magnetospheric cavity expands and can drive outward migration of a planet trapped at the inner edge. They use a one-sided torque operating at the disk edge, based on the results of the studies above. The fact that outward torques operate only on one planet in that model limits the degree to which this effect operates to disrupt a resonance chain, and the resulting simulations still leave too many systems in resonant configurations (Liu & Ormel, 2017). As we will show, the presence of a non-zero torque interior to the location of the torque reversal results in multiple planets migrating outwards and provides a more extensive dynamical perturbation to the system.
@@ 2 Dynamical Model
We wish to model the resonant dynamics of two and three planet systems, under the influence of the most important resonant interactions. We choose to focus on the first order resonances $(q+1)/q$ where $q=1$ and $q=2$. We will also include the effects of torques from the protoplanetary disk on both semi-major axis evolution and planetary eccentricity damping. This is done by including, in the equations, a linear decay term characterised by a timescale, as has been done by several prior authors – (Terquem & Papaloizou, 2019; Choksi & Chiang, 2020). One difference in our case is that our torque is based on the model of Yu et al. (2023), which features a reversal of the torque direction in the inner disk, due to the penetration of the stellar magnetic field into the disk. We must therefore allow for the timescale for semi-major axis evolution ($\tau_{a}$) to become negative, in order to allow for outward migration. The timescale for eccentricity damping is always positive. The equations featuring three planets and two resonances are sufficiently lengthy that we defer their explicit expression to Appendix A.
@@ 2.1 Torque Model
The dynamical evolution of the system is driven by the torques exerted by the protoplanetary disk. We determine a functional form for the torque based on a fit to the numerical disk simulations discussed in Yu et al. (2023). For the inward migration, the characteristic timescale for semi-major axis evolution of planet $i$ is
$\tau_{a,i}$ $=$ $0.36{\rm Myr}\left(\frac{a_{i}}{0.1AU}\right)^{1.38}\left(\frac{M_{i}}{10M_{\oplus}}\right)^{-1}$ (1)
@@
where $a_{i}$ is the semi-major axis, $M_{i}$ is the planetary mass, and $\dot{M}$ is the rate of mass accretion through the disk (assuming a central star of $1M_{\odot}$). The time dependance of $\tau_{i}$ is driven by the evolution of the accretion rate, which is
@@
This is chosen to represent a disk that decays exponentially with time $t$ as it accretes onto the star but then cuts off more rapidly once photoevaporation losses detach the inner disk from mass resupply by the outer disk (Alexander et al., 2014).
Once a planet has crossed interior to the torque reversal line $a_{R}$ – located at the disk surface density maximum (Yu et al., 2023), the migration torque is now directed outwards. We describe this using a negative timescale. The functional form is again determined by fitting to the simulations of Yu et al. (2023), and is expressed as a function of position relative to $a_{R}$, $z=\ln a_{i}-a_{R}$. This amounts to
$\left|\tau_{a,i}\right|$ $=$ $6.66\times 10^{-18}{\rm Myr}\left(\frac{0.1AU}{a_{i}}\right)^{1/2}\left(\frac{10M_{\oplus}}{M_{i}}\right)$
@@
in the case of our default model  [^1] We set the maximum value of $a_{R}=1$AU. The exact value is not important as the migration freezes out long before this occurs.. Indeed, if the migration time inward or outward is shorter than the timescale on which $a_{R}$ is evolving, then the planet will track the torque reversal location, making the relevant characteristic timescale $\tau_{R}=dt/d\ln a_{R}$.
@@
For example, if the outer planet of the pair reaches the torque reversal location, then the pair of planets will both be migrating outwards. The question of whether they converge or diverge will rest on whether the outward migration of the inner planet is faster or slower than the rate at which $a_{R}$ recedes because the outer planet is locked to $a_{R}$ by the fact that the torques on either side of this transition drive the planet towards it. Two examples of this evolution are shown in Figure 1. The solid curve shows the characteristic timescale $\left|\tau_{R}\right|$ associated with the outward movement of the torque reversal location as $\dot{M}$ decreases. We also show two cases of $\left|\tau_{a,i}\right|$ interior to this. The dotted line shows the rate of outward migration for a $6M_{\oplus}$ planet that is the interior member of a pair locked in the 2:1 mean motion resonance, in which the outer member is located at $a_{R}$. At early times, $\left|\tau_{a,i}\right|<\left|\tau_{R}\right|$, so that the planet is able to keep up with the outward migration of the outer member of the pair (which is tied to the torque reversal location $a_{R}$). Thus, the pair is converging and remains trapped in resonance, even during the outward migration. Figure 1 shows that $\left|\tau_{a,i}\right|$ increases with time, while the solid curve ($\left|\tau_{R}\right|$) decreases with time. Therefore, when these two curves cross, the inner planet falls behind and the pair starts to diverge, falling out of resonance.
@@
Figure 1: The solid curve shows the characteristic timescale for the evolution of $a_{R}$ as a function of time. We see that this is initially much longer than the lifetime of the disk, but shortens as time increases, i.e. the torque reversal moves out faster as $\dot{M}$ drops. The dotted curve represents $\left|\tau_{i}\right|$ at a location $0.63a_{R}$, assuming a mass $6M_{\oplus}$. This represents the rate of outward migration for a planet of that mass, trapped in a 2:1 resonance with a planet located at the torque reversal location. The dashed line represents the same but for a $3M_{\oplus}$ planet located at $0.763a_{R}$, i.e. in a 3:2 resonance. Both timescales are shorter than $\left|\tau_{R}\right|$ until about 1.2 Myr. After this, the planets cannot keep up with $a_{R}$ and will freeze out and decouple from the disk. The long dashed line indicates the timescale for inward migration, evaluated at $a_{R}$. This is shorter than the timescales for the inner planet, indicating that the outer member of the pair will freeze out at a later time.
@@
The dashed line shows the case of a $3M_{\oplus}$ planet locked in a 3:2 resonance with an outer member located at $a_{R}$. This shows the same characteristic behaviour, although the divergence starts earlier. The long dashed curve shows the inward migration timescale for a planet located just exterior to $a_{R}$. The fact that this curve falls below the others means that the outer member of the pair remains locked to $a_{R}$ even as the inner member drops out of resonance, although eventually this planet falls behind the retreat of $a_{R}$ as well.
@@
In the following sections, we calculate the evolution of two and three planet systems evolving according to these torques. In addition to the decay of the semi-major axis, these torques also damp eccentricities. Resonant interactions between neighbouring planets lead to a coupled evolution, including eccentricity excitation and resonant lock. The equations that describe this evolution are given in Appendix A.
@@ 3 Converging versus Diverging
where we have chosen the q=1 resonance for this example and the various quantities are defined using the conventions given in appendix A. These conditions also imply specific values for the equilibrium eccentricities. The full expressions are given in Terquem & Papaloizou (2019), but the most important element is that
@@
Thus, when the innermost planet first crosses the torque reversal location, the value of $t_{a,1}$ will change sign. This will cause the pair to converge even more strongly, so that there will still be an equilibrium, although the values of the equilibrium eccentricities will change. This can occur even if the resonant lock does not change. However, once the second planet crosses the torque reversal location, then both planets are moving outwards and the question of whether they are converging or diverging is once again a matter of their relative motion. The equilibrium will fail if the quantity in equation (5) becomes negative – thereby breaking the resonant lock. The system evolution retains the same character if one of the planets is tracking the location of the torque reversal – the relative $t_{a}$ is then simply dictated by $\tau_{R}$ – the evolution of the torque reversal.
@@ 4 Two Planet Systems
Figure 2 illustrates the basic evolution for a two planet system, containing two $3M_{\oplus}$ planets. The initial semi-major axes are 0.076 AU and 0.1 AU, with initial eccentricities of 0.01. The initial mean longitudes and longitudes of periastron were chosen randomly and the relevant resonant angles constructed therefrom. The equations in Appendix A are integrated using the adaptive stepsize integrator odeint from Press et al. (1992). The black curve in the upper left plot illustrates the evolution of the semi-major axis of the inner planet, and the red curve the evolution of the outer planet. The dotted line indicates the evolution of the location of the torque reversal location $a_{R}$. The upper right panel shows the period ratio of the two planets as a function of time. In the lower left, we show the evolution of the planetary eccentricities, and, in the lower right, the panel shows how the resonant angle for the inner and outer 3:2 resonances change with the planetary and disk evolution.
@@
Figure 2: The upper left panel shows the evolution of semi-major axis for a pair of $3M_{\oplus}$ planets migrating due to the gaseous torque model described in the text. The dotted curve shows how the location of the torque reversal evolves with time. Important transitions in the dynamical state of the system are labelled as A–E, and are shown as vertical dashed lines in the other three panels. In the upper right, we see the evolution of the period ratio for the pair. In the lower left we show the evolution of the eccentricity for each planet (red curve indicating the outer planet and black curve the inner planet). In the bottom right, we see the evolution of the cosines of the resonant angles $\phi_{3:2}$, namely $3\lambda_{2}-2\lambda_{1}-\bar{\omega}_{1}$ (black) and $3\lambda_{2}-2\lambda_{1}-\bar{\omega}_{2}$ (red).
@@
The first incident of note – denoted point A – is when the inner planet reaches the torque reversal location $a_{R}$. At this point the inward migration of the inner planet stops and the planet semi-major axis starts to track the evolution of $a_{R}$, which is slowly moving outwards. Up until this point, the period ratio is increasing (the planets are diverging) because the migration time decreases as planets get closer to the star. However, with the halting of the inner planet migration at A, the planets begin to converge because the outer planet, shown as the red curve, continues to migrate inwards, and the two planets enter the 3:2 resonance at B. The net relative motion is convergent and so the pair become locked in a resonant configuration. As a result, the eccentricities rise from almost zero to finite values dictated by the equilibrium defined by equation (5).
@@
The torques on the outer planet now drive the pair inwards, and the inner planet is pushed interior to the torque reversal. Eventually the outer planet reaches the torque reversal (at time denoted as C). The planets remain in equilibrium (because the torques on the inner planet are pushing it outwards, so the pair is still converging) but the equilibrium eccentricities change because the sign of the torques change which changes the equilibrium condition in equation (6). After this, the pair starts to migrate outwards but remains in resonance as long as the rate of outward migration of the inner planet is faster than the evolution of the torque reversal. This continues for a long time, but eventually, as the torques weaken and the outward motion of $a_{R}$ accelerates, the pair starts to diverge, at epoch D. The libration in the 3:2 resonance begins to increase, as does the period ratio, and the eccentricities start to damp away as their resonant excitation weakens. Eventually the migration freezes out completely (at E), leaving the system still in resonance with a finite libration amplitude and an offset relative to strict commensurability.
@@
This is the fundamental result that emerges from these calculations – in the disk models of Yu et al. (2023), planets are not trapped at the true inner edge of the disk, but rather at the location of the torque reversal. As the disk evolves and the surface density drops, this location moves outwards, resulting in an outward migration of the planets at, and interior to, the torque reversal. In many cases, this leads to divergent migration, thereby breaking the resonant lock and causing divergence of the period ratios away from the initial commensurabilities.
@@
The planetary masses also play a role in determining the outcome of the evolution, because more massive planets remain coupled to the gas disk for longer, and lower mass planets freeze out earlier. To illustrate this, Figure 3 shows the case of a two planet system wherein the outer planet ($6M_{\oplus}$) is more massive than the inner one ($3M_{\oplus}$). The initial semi-major axes, eccentricities and longitudes are the same as in Figure 2 and the initial evolution is qualitatively similar, except that the stronger torques on the outer planet means that the inner planet barely halts at the torque reversal before being captured in resonance and pushed further inwards. The eventual divergence from the 3:2 resonance (at D) is also qualitatively similar. However, the outer planet remains coupled to the gas disk for longer by virtue of its larger mass and so migrates far enough that the system passes through the 2:1 resonance (shown at D). This resonant encounter is now divergent, so there is no capture (the system leaves the 2:1 resonance at E) but there is a transient excitation of the eccentricities that occurs late enough to leave small non-zero eccentricities ($\sim 0.01$) after the system has frozen out (at F).
@@
Figure 3: The upper left panel shows the evolution of semi-major axis for a pair of migrating planets where the outer planet is $6M_{\oplus}$ and the inner planet is $3M_{\oplus}$. The dotted curve shows how $a_{R}$ evolves with time. Important transitions in the dynamical state of the system are labelled as A–F, and are shown as vertical dashed lines in the other three panels. In the upper right, we see the evolution of the period ratio for the pair. In the lower left we show the evolution of the eccentricity for each planet (red curve indicating the outer planet and black curve the inner planet). In the bottom right, we have two panels to track the evolution of the cosines of the resonant angles for the two important first order resonances that are encountered. The one designated as $\phi_{3:2}$ shows the angles $3\lambda_{2}-2\lambda_{1}-\bar{\omega}_{1}$ (black) and $3\lambda_{2}-2\lambda_{1}-\bar{\omega}_{2}$ (red). The one designated as $\phi_{2:1}$ shows the evolution of the two resonant angles $2\lambda_{2}-\lambda_{1}-\bar{\omega}_{1}$ (black) and $2\lambda_{2}-\lambda_{1}-\bar{\omega}_{2}$ (red).
@@
The overall amount of migration is driven largely by the planetary masses, as they determine how long the planet remains coupled to the disk and how rapidly the planets migrate with respect to one another. Figure 4 shows the effect of both planetary mass ratio and total mass in the overall evolution of a range of planet pairs. Each pair starts with the same semi-major axes and eccentricities as in Figure 2, while the initial mean longitudes and longitudes of periastron are chosen randomly. We see that low mass ratios ($M_{2}/M_{1}<0.6$) leave pairs trapped in the original 3:2 resonance, whereas larger mass ratios lead to divergence and a wide range of final period ratios, including some in the range of the wider 2:1 resonance. This is a consequence of the fact that the more massive planets remain coupled to the gas for longer. Thus, if the inner planet is the more massive one, the pair will remain in the convergent state for the entirety of the disk evolution. However, if the outer planet is the more massive one, then the inner planet will freeze out first and fall behind the outer planet’s migration. This leads to a diverging configuration and a final period ratio that increases as $M_{2}/M_{1}$ does. The degree of divergence also depends on the total planet mass, with more massive planetary systems leading to large final mass ratios (because they remain coupled to the disk for longer).
@@
Figure 4: The lower panel shows the final period ratio of a planetary pair as a function of the mass ratio $M_{2}/M_{1}$. The results are color coded according to the final value, with black signifying $P_{2}/P_{1}<1.55$, red (1.55–1.95), magenta (1.95–2.05), cyan (2.05–3) and green for $P_{2}/P_{1}>3$. In the upper panel, these colors indicate the final states of systems as a function of mass ratio and total system mass $M_{1}+M_{2}$. The dotted lines in the lower panel indicate the first order commensurabilities (1+q)/q for q=2 and q=1.